# Competitive and Revenue-Optimal Pricing with Budgets

### v2

@@ Abstract
In markets with budget-constrained buyers, competitive equilibria need not be efficient in the utilitarian sense, or maximise the seller’s revenue. We consider a setting with multiple divisible goods. Firstly, we show that competitive equilibrium outcomes, and only those, are constrained utilitarian efficient, a notion of utilitarian efficiency that respects buyers’ demands and budgets. Secondly, we establish that, when buyers have linear valuations, competitive equilibrium prices are unique and revenue-optimal for a zero-cost seller.
@@ 1 Introduction
We consider the notion of constrained social welfare, i.e., social welfare among all allocations and prices [^2] Any feasible allocation is inherently tied to prices in our market, as feasibility must respect the buyers’ budgets. that respect the buyers’ demands, budgets, and market supply. Our first main result shows that, when buyers have quasi-linear preferences, constrained social welfare is maximised in a competitive equilibrium, and in a competitive equilibrium only (Theorem 1). [^3] This theorem continues to hold if the seller has quasi-linear utility with non-zero costs. We also say that a competitive equilibrium is constrained utilitarian efficient, or simply constrained efficient. As is standard, a competitive equilibrium is Pareto efficient, but not necessarily utilitarian efficient (see also Section 2).
@@
As our second main result, we show that, when buyers have linear valuations, the unique competitive equilibrium also maximises the seller’s revenue in a price-only mechanism (Theorem 2). The unique market-clearing prices are buyer-optimal among all revenue-maximising prices, as they maximise the quantities allocated to each buyer. However, there exist revenue-maximising outcomes that are not the competitive equilibrium, and therefore are not constrained utilitarian efficient. We prove our second main result by studying the feasible region, defined as the set of prices at which, for every good, either the market clears (respecting buyers’ budgets and demands) or there is excess supply. We show that this non-convex region has elementwise-smallest prices, and that these prices clear the market as well as well maximising the seller’s revenue.
@@
The notion of constrained utilitarian efficiency provides a compelling and practical benchmark for social optimality in markets with budgets, where competitive equilibrium allocations need not be classically utilitarian efficient. While competitive equilibria in our market remain Pareto efficient, welfare considerations require the stronger criterion of utilitarian efficiency. Studying the constrained social optimum is especially important in settings in which it would be hard to accept breaking or circumventing budget constraints for the sake of efficiency. This is the case in our two motivating examples, digital advertising, and financial asset exchanges. In an ad auction, although a small business may derive a high value from an ad placement, only their ability to pay can be relevant to a for-profit digital platform. Similarly, in the exchange of financial assets, the budget of a buyer represents the nominal amount limit of their asset that is to be exchanged, and thus the seller would not want to grant an allocation of substitute assets exceeding this limit. [^4] In general, the analysis of social welfare that transcends buyers’ budget constraints may be more relevant when these budget constraints are linked to individuals with unequal endowments or socially indispensable businesses with unequal access to capital markets.
@@
Our market has been studied as a simple model of many ad auctions as they occur on digital market platforms (Conitzer et al., 2022), and is also called a mixed Fisher or quasi-Fisher market (Chen et al., 2007; Murray et al., 2020). When businesses compete for digital advertising space, the decision of which publisher (product) to bid for is non-trivial. It is intuitive to choose an advertising budget and state demand in terms of “limit market prices” for multiple, distinct products. The seller or platform assigns to each buyer those products that yield the highest value for money to them. Our second result suggests that the platform can set prices that are both revenue maximizing and socially optimal within the feasibility constraints of advertiser budgets. However, it does not imply that the seller always sets socially optimal prices.
@@ 2 Examples: efficiency, constrained efficiency, and revenue
We illustrate the objectives of maximizing social welfare, constrained social welfare, and the seller’s revenue. Example 1 shows that competitive equilibrium with budget constraints, which is Pareto efficient, is also constrained utilitarian efficient, but not necessarily fully utilitarian efficient. In Example 2, we show that with diminishing marginal values, the seller’s revenue is not maximised in a constrained-efficient allocation, and Example 3 illustrates, in a setting with two goods and constant marginal values, the coincidence of competitive equilibrium and revenue-optimality when buyers have constant marginal values. Note that allocations are non-negative throughout the paper.
@@
Maximising revenue with a price-only mechanism, the seller sets a price in anticipation of the buyer’s demand. The buyer demands the quantity $x$ at price $p$ that maximises $v(x)-pb$ subject to not exceeding the budget, $px\leq\beta$. Thus, the seller’s revenue is maximised at price $p$ and allocation $x\leq s$ that maximise $px$ and the buyer demands $x$ at $p$. Social welfare, on the other hand, is highest at price $p$ and an allocation $x\leq s$ that maximises $v(x)$ among all $(p,x)$ for which the buyer demands $x$ at $p$.
@@ 4 Constrained efficiency
Our first main result establishes that competitive equilibrium maximises constrained efficiency, and any feasible outcome maximising constrained efficiency constitutes a competitive equilibrium. [^10] These results are reminiscent of the fundamental welfare theorems. However, we consider utilitarian efficiency instead of Pareto efficiency. The standard welfare theorems are known to hold in the Arrow-Debreu exchange economy, which embeds our market.
Theorem 1.
A market outcome is a competitive equilibrium if and only if it is constrained efficient.
@@
We note that Theorem 1 continues to hold if the seller has non-zero costs. Costs for sold supply or, equivalently, utility for retained supply may be relevant in the application of financial asset exchanges. In most cases, leftover assets have value to the seller, or the issuance of financial instruments comes at a cost.
@@ 5 The revenue-welfare coincidence
We now assume that each buyer has a linear valuation. In this linear market setting, we establish that competitive equilibrium prices are unique and maximise not just constrained social welfare (Theorem 1) but also revenue. However, while Theorem 1 tells us that constrained social welfare is maximised only at these prices, maximum revenue may also be attained at higher prices, at which the seller only sells a subset of her supply.
@@ 5.2 Maximising revenue and welfare
We show that, for any $\bm{p}\leq\bm{q}$, the maximum revenue obtainable at $\bm{p}$ is weakly greater than the revenue obtainable at $\bm{q}$. It then follows immediately that revenue is maximised at $\bm{p}^{*}$.
@@
Theorem 2 now follows immediately from Propositions 6 and 5.

### v3

@@ Abstract
In markets with budget-constrained buyers, competitive equilibria need not be efficient in the utilitarian sense, or maximise the seller’s revenue. We consider a setting with multiple divisible goods. Competitive equilibrium outcomes, and only those, are constrained utilitarian efficient, a notion of utilitarian efficiency that respects buyers’ demands and budgets. Our main contribution establishes that, when buyers have linear valuations, competitive equilibrium prices are unique and revenue-optimal for a zero-cost seller.
@@ 1 Introduction
We consider the notion of constrained social welfare, i.e., social welfare among all allocations and prices [^2] Any feasible allocation is inherently tied to prices in our market, as feasibility must respect the buyers’ budgets. that respect the buyers’ demands, budgets, and market supply. When buyers have quasi-linear preferences, constrained social welfare is maximised in a competitive equilibrium, and in a competitive equilibrium only (Proposition 4). [^3] This theorem continues to hold if the seller has quasi-linear utility with non-zero costs. We also say that a competitive equilibrium is constrained utilitarian efficient, or simply constrained efficient. As is standard, a competitive equilibrium is Pareto efficient, but not necessarily utilitarian efficient (see also Section 2).
@@
Our main result establishes that when buyers have linear valuations, the unique competitive equilibrium also maximises the seller’s revenue in a price-only mechanism (Theorem 1). The unique market-clearing prices are buyer-optimal among all revenue-maximising prices, as they maximise the quantities allocated to each buyer. However, there exist revenue-maximising outcomes that are not the competitive equilibrium, and therefore are not constrained utilitarian efficient. We prove our result by studying the feasible region, defined as the set of prices at which, for every good, either the market clears (respecting buyers’ budgets and demands) or there is excess supply. We show that this non-convex region has elementwise-smallest prices, and that these prices clear the market as well as maximise the seller’s revenue.
@@
The notion of constrained utilitarian efficiency provides a compelling and practical benchmark for social optimality in markets with budgets, where competitive equilibrium allocations need not be classically utilitarian efficient. While competitive equilibria in our market remain Pareto efficient, welfare considerations require the stronger criterion of utilitarian efficiency. Studying the constrained social optimum is especially important in settings in which it would be hard to accept breaking or circumventing budget constraints for the sake of efficiency. This is the case in our two motivating examples, digital advertising, and financial asset exchanges. In an ad auction, although a small business may derive a high value from an ad placement, only their ability to pay can be relevant to a for-profit digital platform. Similarly, in the exchange of financial assets, the budget of a buyer represents the limit on the nominal value of their asset that is to be exchanged, and thus the seller would not want to grant an allocation of substitute assets exceeding this limit. [^4] In general, the analysis of social welfare that transcends buyers’ budget constraints may be more relevant when these budget constraints are linked to individuals with unequal endowments or socially indispensable businesses with unequal access to capital markets.
@@
Our market has been studied as a simple model of many ad auctions as they occur on digital market platforms (Conitzer et al., 2022), and is also called a mixed Fisher or quasi-Fisher market (Chen et al., 2007; Murray et al., 2020). When businesses compete for digital advertising space, the decision of which publisher (product) to bid for is non-trivial. It is intuitive to choose an advertising budget and state demand in terms of “limit market prices” for multiple, distinct products. The seller or platform assigns to each buyer those products that yield the highest value for money to them. Our main result suggests that the platform can set prices that are both revenue maximizing and socially optimal within the feasibility constraints of advertiser budgets. However, it does not imply that the seller always sets socially optimal prices.
@@ 2 Examples: Efficiency, Constrained Efficiency, and Revenue
We illustrate the objectives of maximizing social welfare, constrained social welfare, and the seller’s revenue. Example 1 shows that competitive equilibrium with budget constraints, which is Pareto efficient, is also constrained utilitarian efficient, but not necessarily fully utilitarian efficient. In Example 2, we show that with diminishing marginal values, the seller’s revenue is not maximised in a constrained-efficient allocation. Example 3 illustrates our main result (Theorem 1), the coincidence of competitive equilibrium and revenue-optimality, in a setting with two goods and constant marginal values. Note that allocations are non-negative throughout the paper.
@@
Maximising revenue with a price-only mechanism, the seller sets a price in anticipation of the buyer’s demand. The buyer demands the quantity $x$ at price $p$ that maximises $v(x)-px$ subject to not exceeding the budget, $px\leq\beta$. Thus, the seller’s revenue is maximised at price $p$ and allocation $x\leq s$ that maximise $px$ and the buyer demands $x$ at $p$. Social welfare, on the other hand, is highest at price $p$ and an allocation $x\leq s$ that maximises $v(x)$ among all $(p,x)$ for which the buyer demands $x$ at $p$.
@@ 4 Constrained Efficiency
First, we establish that competitive equilibrium maximises constrained efficiency, and any feasible outcome maximising constrained efficiency constitutes a competitive equilibrium. [^10] This is reminiscent of the fundamental welfare theorems, but we consider utilitarian efficiency instead of Pareto efficiency. The standard welfare theorems are known to hold in the Arrow-Debreu exchange economy, which embeds our market. This classic result is well-known for markets with indivisible goods without budgets (see, e.g., Sun and Yang (2014)) and we extend this to our market in which outcomes must respect budgets, demand and supply.
Proposition 4.
A market outcome is a competitive equilibrium iff it is constrained efficient.
@@
We note that Proposition 4 continues to hold if the seller has non-zero costs. Costs for sold supply or, equivalently, utility for retained supply may be relevant in the application of financial asset exchanges. In most cases, leftover assets have value to the seller, or the issuance of financial instruments comes at a cost.
@@ 5 The Revenue-Welfare Coincidence
We now assume that each buyer has a linear valuation. In this linear market setting, we establish that competitive equilibrium prices are unique and maximise not just constrained social welfare (Proposition 4) but also revenue. However, while Proposition 4 tells us that constrained social welfare is maximised only at these prices, maximum revenue may also be attained at higher prices, at which the seller only sells a subset of her supply.
@@ 5.2 Maximising Revenue and Welfare
We show that, for any feasible $\bm{p}\leq\bm{q}$, the maximum revenue obtainable at $\bm{p}$ is weakly greater than the revenue obtainable at $\bm{q}$. It then follows immediately that revenue is maximised at $\bm{p}^{*}$.
@@
Theorem 1 now follows immediately from Propositions 7 and 6.